OpenStudy (anonymous):

What's the second derivative of f(x)= sec^2 (pix)?

OpenStudy (anonymous):

cos^2

OpenStudy (anonymous):

^incorrect.

OpenStudy (anonymous):

i thnk so

OpenStudy (anonymous):

The answer would be: \[\frac{d}{dx} \left( \frac{d}{dx}(\sec^2(\pi x)\right)=\frac{d}{dx} \left( 2 \pi(\sec^2( \pi x)\tan(\pi x) \right)\] From the product rule. If you look CAREFULLY you see that if we differentiate whats left we do the product rule, one of them involves the derivative of sec^2(pi x) which we JUST FOUND, so use that result to get: \[4 \pi^2 \sec^2(\pi x)\tan^2(\pi x)+2 \pi^2 \sec^4(\pi x)\]

OpenStudy (anonymous):

Not cos^2